### v3

# MathScape: Evaluating MLLMs in multimodal Math Scenarios through a Hierarchical Benchmark

Abstract
With the development of Multimodal Large Language Models (MLLMs), the evaluation of multimodal models in the context of mathematical problems has become a valuable research field. Multimodal visual-textual mathematical reasoning serves as a critical indicator for evaluating the comprehension and complex multi-step quantitative reasoning abilities of MLLMs. However, previous multimodal math benchmarks have not sufficiently integrated visual and textual information. To address this gap, we proposed MathScape, a new benchmark that emphasizes the understanding and application of combined visual and textual information. MathScape is designed to evaluate photo-based math problem scenarios, assessing the theoretical understanding and application ability of MLLMs through a categorical hierarchical approach. We conduct a multi-dimensional evaluation on 11 advanced MLLMs, revealing that our benchmark is challenging even for the most sophisticated models. By analyzing the evaluation results, we identify the limitations of MLLMs, offering valuable insights for enhancing model performance. The code is made available https://github.com/PKU-Baichuan-MLSystemLab/MathScape.
1 Introduction
In recent years, there have been advancements in large language models (LLMs) (OpenAI 2023a; Touvron et al. 2023) and MLLMs (Zhao et al. 2023; Wu et al. 2023; Bai et al. 2024). They have shown strong understanding ability among different modalities (Liu et al. 2023b; Bai et al. 2023b).
Figure 1: MathScape offers a comprehensive collection of math problems from primary school to high school. The problems range in difficulty from easy to difficult, catering to various levels of evaluation.
Among Multimodal Large Language Models (MLLMs), Vision Language Large Models (VLLMs) have demonstrated competitive performance in traditional multimodal tasks, including image classification (Chen et al. 2024), image understanding (Li et al. 2023b, c), and image captioning (Bai et al. 2023b). Furthermore, their advanced language understanding capabilities contribute to strong performance in text-rich tasks, such as visual question answering (Liu et al. 2023b, a) and image-text retrieval (Chen et al. 2024). Recently, VLLMs have also shown significant progress in solving mathematical problems. Therefore, comprehensive benchmarks are essential to evaluate the mathematical abilities of VLLMs. Although several benchmarks, such as MATH-V (Wang et al. 2024a), MathVerse (Zhang et al. 2024), and MathVista (Lu et al. 2023b), have been developed to assess the mathematical capabilities of VLLMs. They primarily focus on a combination of text math problems and image figures. Also, they only use simple metrics and lack effective evaluation for complex or extended responses. Consequently, they face two key challenges:
C1. Insufficient Real-World Data. In previous datasets like MATH-V (Wang et al. 2024a), MathVerse (Zhang et al. 2024), and MathVista (Lu et al. 2023b), the mathematical description was typically provided as text input, while the image contained only figures. This approach doesn’t align well with real-world scenarios, where both the mathematical description and figures are captured together in a single image.
C2. Absence of Effective Evaluation Metrics. In previous datasets (Wang et al. 2024a; Zhang et al. 2024; Lu et al. 2023b), the evaluation was limited to short answers, lacking the ability to assess long-form responses.
To address these issues, we implement a three-step pipeline for constructing a real-world math image dataset. As illustrated in Figure 3, the process begins by converting math documents into images, as shown in Figure 2. Next, we capture photos and screenshots to build the dataset. Finally, we perform a thorough review and knowledge classification to ensure the dataset’s high quality. For evaluation, we design a two-step pipeline specifically for assessing longer math problems. First, we use LLMs to extract answers for each subproblem. Then, we employ LLMs as evaluators to assess the correctness of each solution. With the data construction and evaluation pipeline, we constructed MathScape, a new multimodal dataset that combines photos of real-world math problems with their correct answers.
Figure 2: An problem example of MathScape. Examples in MathScape are represented by images taken by humans, ensuring a more realistic scenario. Each example will contain a correct answer.
The core contributions are summarized as follows:
  • New Perspective: To the best of our knowledge, we are the first to construct images that combine both figures and mathematical text descriptions, closely mirroring real-world scenarios.
  • New Method: We propose a novel three-step dataset construction pipeline, as illustrated in Figure 3. Additionally, we introduce a new two-step evaluation method specifically designed for assessing long answers.
  • New Benchmark: We present MathScape, a new multimodal mathematical dataset that spans various difficulty levels, question types, and knowledge areas, providing a comprehensive tool to evaluate the mathematical capabilities of MLLMs. Moreover, MathScape is entirely original, consisting of previously unreleased multimodal mathematical data.
2 Related Work
In the field of MLLMs, the benchmark for multimodal mathematical reasoning capability represents a significant and novel research direction. Mathematical reasoning is a crucial indicator for evaluating the ability of LLMs to perform complex, multi-step reasoning and quantitative analysis within visual contexts. Below, we highlight some relevant work and the latest developments in this area.
2.1 Benchmark for Mathematical Evaluation
Recent research has seen significant advancements in mathematical reasoning benchmarks aimed at evaluating mathematical abilities. In this summary, we review both pure text and multimodal math benchmarks.
Pure Text Benchmarks
GSM8K (Cobbe et al. 2021) is a dataset from OpenAI that includes 8.5K high-quality elementary school math word problems, each requiring 2 to 8 steps to solve. These problems primarily involve basic arithmetic operations such as addition, subtraction, multiplication, and division. MATH (Hendrycks et al. 2021) offers a dataset of 12,500 problems sourced from high school math competitions. SuperCLUE-Math (Xu et al. 2024a) is a Chinese benchmark for multi-step reasoning in mathematics, containing over 2,000 problems that require multi-step reasoning and offer natural language solutions. MathBench (Liu et al. 2024b) includes 3,709 math problems ranging from basic arithmetic to college-level questions, covering multiple difficulty levels.
All these benchmarks focus exclusively on text-based mathematical tasks. They are designed to evaluate the mathematical capabilities of LLMs through specialized problem sets.
Multimodal Benchmarks
With the rapid advancement of MLLMs, several high-quality benchmarks have emerged to evaluate mathematical problem-solving in visual contexts. MathVista (Lu et al. 2023b) focuses on visual math QA tasks, assessing model performance across various math domains, such as arithmetic and algebra, using visual scenarios. MATH-V (Wang et al. 2024a) is another benchmark that targets multimodal mathematical understanding, with questions primarily sourced from math competitions. MathVerse (Zhang et al. 2024) evaluates MLLMs’ comprehension of visual diagrams using CoT (Chain of Thought) strategies on 2,612 multimodal math problems. CMMU (He et al. 2024) is a large-scale Chinese benchmark for multi-disciplinary, multimodal understanding, featuring questions from college exams and textbooks.
Compared to these existing multimodal mathematical benchmarks, which often have limitations in question length, complexity, and openness to model answers, our MathScape benchmark is designed to be longer and more open-ended.
2.2 MLLMs for Mathematics
Commonly Used VLLMs
The integration of visual knowledge into LLMs has become a pivotal area of research due to the rapid advancements in LLMs. VLLMs combine vision information from vision encoders with LLMs, thus enabling these models to process and interpret visual inputs for various visual tasks  (Liu et al. 2023c; Zhang et al. 2022; Li et al. 2022b) with enhanced accuracy and efficiency. Pioneering frameworks like CLIP  (Radford et al. 2021) leverage contrastive learning on expansive image-caption datasets to align modalities, forming the groundwork for cross-modal comprehension. Various adapters  (Liu et al. 2023b, a; Li et al. 2023b, 2022a; Jian, Gao, and Vosoughi 2023; Lu et al. 2023a) are introduced to further integrate different modalities. For example, LLaVA  (Liu et al. 2023b, a) employs a straightforward MLP to inject the vision information into LLMs. Whereas more complex implementations like the Q-Former in BLIP  (Li et al. 2022a, 2023b) utilize cross-attention to enhance modality integration.
Recent studies  (Wang et al. 2024b; Chen et al. 2023; Liu et al. 2023b, a; Li et al. 2023a) aims to boost VLLM performance by focusing on the quality of both pre-training and fine-tuning datasets. Models like LLaVA  (Liu et al. 2023b, a) and ShareGPT4V  (Chen et al. 2023) have shown remarkable advancements in understanding and following complex instructions through instruction tuning.
VLLMs Designed for Math Problems
MAmmoTH (Yue et al. 2023) InternLM-Math (Ying et al. 2024), and ChatGLM-Math (Xu et al. 2024b) are multimodal models specifically tailored for dealing with mathematical questions, incorporating both textual and visual components in their problem design to enhance their ability to handle complex mathematical tasks.
3 Methodology
We begin by introducing the construction pipeline of MathScape in Section 3.1. Next, we present the multidimensional evaluation approach in Section 3.2. In Section 3.3, we detail the two-step answer evaluation method. Finally, we summarize the dataset statistics in Section 3.4.
Figure 3: MathScape process pipeline.
3.1 Construction of MathScape
Data Preparation
The data preparation module consists of three steps, as shown in Figure 3(a). First, we collected a large number of mathematics questions from elementary, junior high, and senior high school exams and homework as the evaluation sample. We gathered a total of 1,325 image mathematics questions. Next, the question documents were converted to PDF format using Pandoc and subsequently transformed into images for further use.
Data Annotation
As illustrated in Figure 3(b), the images are then transformed to closely align with real-world scenarios by capturing photos of printed images and screen displays.
Data Check and Knowledge Classification
After constructing the dataset, we perform a double-check and knowledge-based classification to ensure its high quality. As illustrated in Figure 3(c), we rigorously review the dataset to ensure that both the textual and graphical inputs are clear and accurate. Once data quality is verified, we categorize the data according to knowledge points.
3.2 Multidimensional Evaluation
To comprehensively evaluate the performance of VLLMs, we designed multiple dimensions to classify and assess their mathematical abilities across various categories. The classification types we used are as follows:
Question Types:
We first categorized the test questions into different types, such as multiple-choice, fill-in-the-blank (Solution), and proof questions, to examine the model’s performance across various question formats.
Knowledge Points:
We also classified the questions based on mathematical knowledge areas, including algebra, geometry, probability, and statistics, to assess the model’s proficiency in different domains of mathematics.
Educational Stages:
Additionally, the questions were divided according to the educational stage—primary school, middle school, and high school—to evaluate the model’s adaptability and accuracy at different levels of education.
3.3 Evaluation Method
We utilize a two-step evaluation process to effectively score long answers.
Answer Segmentation:
As illustrated in Figure 4, we prompt the LLMs to decompose a lengthy answer into multiple sub-answers, each one focusing on a specific aspect of the problem. This segmentation ensures that the complex answer is broken down into manageable components, making it easier to evaluate the correctness and relevance of each part. By isolating sub-problems within the overall solution, we can achieve a more granular analysis of the model’s performance.
Sub-Answer Scoring:
After segmenting the long answer, we employ the prompt depicted in Figure 12 to automatically score each sub-answer individually. This method allows us to evaluate the accuracy of each component independently, ensuring that the final score reflects the model’s ability to handle various aspects of the problem comprehensively. By scoring sub-answers separately, we can identify specific areas where the model excels or struggles, providing deeper insights into its strengths and weaknesses.
Figure 4: Prompts for inference and extracting answers.
Figure 5: Proportion Figure
3.4 Dataset Statistics
In this section, we provide a summary of the statistics for our MathScape dataset. The dataset primarily consists of Chinese image-text problems, along with question labels, attribute information, problem-solving processes, and standard reference answers. Detailed statistics are presented in Figure 5.
As shown in Figure 5(a), our dataset thoughtfully incorporates the characteristics of multimodal image-text questions. A significant portion of the questions are geometric, which often require the integration of images for effective problem-solving. In contrast, topics like equations and inequalities are less represented, aligning more closely with the specific demands of multimodal assessment.
Figure 5(b) illustrates that our dataset primarily includes solution questions and multiple-choice questions, with fewer proof questions. This distribution indicates that our dataset is designed to challenge models with diverse question types, while still reflecting the real-world emphasis on practical problem-solving.
Overall, our dataset contains a total of 1,325 images, providing a robust resource for evaluating the mathematical reasoning capabilities of MLLMs.
Table 1: Accuracy scores comparison of models on different question types
Model Average Choice Solution Proof
Closed-source Models
GPT-4V 34.96 35.75 31.72 28.33
GPT-4-turbo 33.92 29.85 31.58 56.62
Claude-3-Opus 28.79 29.3 20.85 50.00
Gemini-Pro 21.37 12.62 16.16 37.50
Baichuan-VL 30.00 26.38 25.83 45.97
Qwen-VL-Max 27.83 23.97 22.17 34.85
Qwen-VL-Plus 15.60 19.46 12.48 35.19
GLM4V 12.26 11.54 7.31 26.28
Open-source Models
Yi-VL-34B 18.36 19.01 9.98 33.33
DeepSeek-V2 15.66 12.75 10.60 37.69
LLaVA-1.6-7B 12.35 11.31 6.24 13.43
Table 2: Accuracy scores comparison of Models on different knowledge points
Model Algebraic Geometric Equations and Inequality Functions Probability Statistics
Closed-source Models
GPT-4V 39.05 27.90 29.73 34.14 41.31
GPT4-turbo 36.28 29.54 32.50 28.43 37.99
Claude-3-Opus 31.78 22.67 20.83 20.58 36.22
Gemini-Pro 21.13 15.50 15.35 9.57 13.33
Baichuan-VL 30.54 25.98 25.83 26.69 23.67
Qwen-VL-Max 28.71 21.86 28.33 20.86 19.09
Qwen-VL-Plus 16.70 17.07 18.67 16.67 11.46
GLM4V 8.94 12.57 5.13 7.32 10.55
Open-source Models
Yi-VL-34B 16.78 15.84 7.02 9.79 11.44
DeepSeek-V2 12.71 14.87 6.19 10.60 9.61
LLaVA-1.6-7B 9.76 8.58 15.79 3.57 10.77
4 Experiments and Analysis
In this section, we utilize multiple state-of-the-art (SOTA) models and test their performance on the MathScape benchmark.
4.1 Experimental Setups
Models.
In our evaluation of multimodal LLMs, we focused on both open-source and closed-source models that rank among the top performers on major multimodal LLM leaderboards. This included 11 different types of VLLMs, with a particular emphasis on analyzing the results and performance of the leading models.
  • Closed-source models: GPT4 (OpenAI 2023b), GeminiPro (Reid et al. 2024), Claude-3-Opus, Baichuan-VL (Yang et al. 2023), Qwen-Max (Bai et al. 2023a), Qwn-Plus (Bai et al. 2023a), GLM4V.
  • Open-source models: Deepseek-VL(Lu et al. 2024), LLaVA(Liu et al. 2024a), Yi(Young et al. 2024).
Settings.
We conduct all model inferences in a zero-shot setting, using the same configuration for each official model. Instead of the Chain of Thought (CoT) technique, we use a custom prompt to guide the model in producing the problem-solving process and final answer, as shown in Figure 4. The settings include a max token limit of 2048, top-k of 5, a temperature of 0.3, and a repetition penalty of 1.05. All experiments are run on NVIDIA H100 GPUs.
Table 3: Comparison of Models on different knowledge stages (E: Easy, M: Medium, D: Difficult, Avg: Average Score)
Model Elementary Middle High
avg E M D avg E M D avg E M D
Closed-source Models
GPT-4V 36.04 57.58 38.64 10.71 36.42 40.38 34.95 30.14 28.08 33.26 24.38 22.57
GPT4-turbo 37.71 72.73 38.79 18.33 35.12 37.22 34.51 30.44 26.06 28.65 25.19 18.83
Claude-3-Opus 28.30 33.33 31.10 10.04 31.04 31.29 33.97 12.22 19.17 24.07 16.41 15.15
Gemini-Pro 25.79 48.48 26.91 11.29 17.20 19.19 16.29 15.07 10.22 12.74 8.90 5.03
Baichuan-VL 29.85 35.00 31.45 18.33 29.96 28.94 32.57 21.38 22.33 27.59 17.42 16.01
Qwen-VL-Max 34.82 42.86 36.65 20.45 24.87 25.70 24.96 20.72 16.95 18.97 15.61 14.92
Qwen-VL-Plus 20.49 40.00 21.23 9.20 19.16 21.11 18.83 13.19 11.00 13.94 9.29 5.83
GLM4V 10.32 33.29 9.62 4.29 13.28 17.07 14.85 12.89 7.64 8.73 11.11 4.08
Open-source Models
Yi-VL-34B 14.99 40.00 16.13 3.32 16.38 16.31 17.10 11.67 12.14 11.65 12.96 10.58
DeepSeek-V2 13.74 42.42 13.73 2.87 14.93 14.68 14.47 19.09 10.18 8.29 12.46 7.99
LLaVA-1.6-7B 9.77 35.21 10.82 7.12 10.37 9.79 10.90 9.07 7.57 8.41 6.53 4.54
4.2 Performance of Various Models
In this section, we present the performance of commonly used MLLMs on our benchmark. We analyze the results from the perspectives of Question Types, Knowledge Points, and Educational Stages:
Question Types
As shown in Table 1, GPT-4V and GPT-4-turbo exhibits the highest accuracy across all question types, with an average of 34.96%, followed by GPT-4-Turbo Vision at 33.92%. While Yi-VL-34B and DeepSeek-V2 achieve good performance among open-source models. We can see the performance of closed-source models achieved better performance than open-source models.
The table shows that models generally perform better on proof questions compared to multiple-choice and solution questions. This suggests that the structured format and clear information in proof questions make them easier for models to handle, while solution questions, which require complex, multi-step reasoning, pose more of a challenge.
Knowledge Points
Table 2 shows the answer accuracy of the models in different knowledge points. GPT-4V and GPT-4-turbo consistently outperform other models in areas like algebra, equations and inequalities, functions, and probability and statistics. Most models show balanced performance across different knowledge areas, but there are exceptions, such as LLaVA-1.6, which does well in equations and inequalities but struggles with functions.
Overall, closed-source models are more accurate than open-source ones, with GPT-4V and GPT-4-turbo leading in many categories.
Educational Stages
Table 3 presents the performance of open-source and closed-source models on MathScape at the elementary, middle, and high school levels. At the elementary and middle levels, the models perform similarly. However, when the difficulty increases to the high school level, we observe a significant drop in accuracy. Some models show an extreme decrease in performance between the middle and high school benchmarks. For instance, Gemini-Pro has an average accuracy of 25.79% at the elementary level, but this sharply declines to just 10.22% at the high school level. This suggests that high school-level math poses significant challenges for LLMs.
Overall, our evaluation shows that closed-source models, particularly GPT-4V and GPT-4-turbo, consistently outperform open-source models across various question types, knowledge points, and educational stages. These models demonstrate superior accuracy, especially in structured question types like proof questions and in areas requiring advanced mathematical reasoning, such as algebra and probability. However, as the difficulty level increases, all models experience a decline in accuracy, with the most significant drops occurring between the middle and high school stages. GLM-4V performs particularly poorly at the high school level, highlighting the challenges that remain in achieving consistent performance on difficult math problems.
Figure 6: Stability Analysis: For each problem, the model is tested five times. The numbers 1 to 5 represent the proportion of correct responses.
Figure 7: The variation of accuracy with answer length.
4.3 Stability Results and Analysis
In this subsection, we perform a stability test for GPT4V, Claude-3-Opus, Baichuan-VL, and Qwen-VL-Max. We selected 300 problems and tested each model five times on each problem. The number of correct answers across these attempts was calculated to assess the stability of each model. As shown in Figure 6, none of the models demonstrate high stability—only about 25% of the problems were answered correctly in all five attempts. Therefore, it’s imperative to focus on enhancing the stability and robustness of math MLLMs, as consistent performance across repeated trials is crucial for their practical application in real-world scenarios. This finding also suggests that future research should explore methods to reduce variability in model outputs, ensuring more reliable and trustworthy results.
4.4 Answer Length and Accuracy
Distribution of Answer Lengths
From Figure 7, we observe distinct patterns in the distribution of answer lengths across different models. Notably, GPT-4V and Baichuan-VL tend to generate a larger proportion of shorter answers. As illustrated in Figure 8, it is evident that shorter but accurate answers are more likely to achieve higher scores. This trend highlights the efficiency of models that can deliver concise and precise responses, particularly in scenarios where brevity is valued.
Figure 8: The variation of accuracy with answer length.
Figure 9: Illustration of Case Study 1. The red sign marks the incorrect part of the answer.
Analysis of Answer’s Length
In our evaluation of the MathScape benchmark, we observed that there is no straightforward positive correlation between answer length and accuracy. In fact, as shown in Figure 8, when the length of the answer increases, the accuracy tends to decrease. This result demonstrates the robustness of the MathScape benchmark, ensuring that models cannot simply inflate their scores by producing longer answers. Such a design effectively prevents any biases in answering strategies, ensuring that the benchmark and evaluation method accurately reflects a model’s true ability to understand and solve mathematical problems, rather than gaining an unfair advantage through verbose responses.
5 Challenges and Future Directions
As shown in Section 4, none of the models performed well on our MathScape benchmark. In this section, we discuss the challenges faced by current MLLMs and suggest potential future directions for improving math MLLMs.
5.1 Challenges
In this subsection, we explore the main reasons why models provide incorrect answers to image-text mathematical problems. These errors are mainly due to challenges in understanding and interpreting the information. We can break down these challenges into the following specific reasons:
Unable to Retrieve Information from the Image:
This is one of the most common errors, where models may fail to extract all the relevant information from the image. For instance, when interpreting complex geometric patterns, it’s easy to overlook certain data or conditions, leading to incorrect answers. As shown in Case Study 1 in Figure 9, the model provided an incorrect proof due to its incomplete understanding of the image.
Misunderstanding of Graphic Positioning:
This issue involves the accurate understanding of the spatial layout of graphics. For instance, in geometry problems, errors can occur if the model fails to correctly recognize the lengths or angles of figures. Such mistakes often stem from a lack of deep understanding of graphic properties or insufficient ability to shift perspectives. In Figure 10, Case Study 2, the model incorrectly interprets the distance from point A to 0 as $2\sqrt${2}square-root start_ARG 2 end_ARG.
Insufficient Reasoning Ability:
This issue arises from the limited logical reasoning capabilities of LLMs. Even when the image information is provided correctly, the LLM may still produce incorrect responses. As shown in Case Study 3 in Figure 11, the LLM fails to solve the complex problem correctly and makes errors in the process.
Overall, the challenges for multimodal large models primarily focus on the interpretation of visual information and the inherent reasoning abilities of the models.
5.2 Future Directions for Math MLLMs
MathScape have introduced several challenges for MLLMs, as mentioned in section 5.1. In this section, we summarize future directions for MLLMs.
Stronger LLMs
As discussed in section 5.1, it is evident that LLMs have limited mathematical reasoning abilities. To improve the math problem-solving capabilities of MLLMs, it is crucial to develop stronger LLMs with enhanced mathematical reasoning skills.
Better Pattern Recognition
Improving pattern recognition is essential for enhancing the performance of MLLMs, particularly in tasks involving complex visual information. Current models often struggle with identifying and interpreting intricate patterns in images, such as geometric configurations, charts, and fine-grained visual details. Future research should focus on developing models that can more accurately recognize and differentiate patterns, especially when they are complex.
6 Conclusion
Recently, MLLMs have emerged as powerful models for answering questions across multiple domains. However, comprehensive benchmarks that reflect real-world scenarios are needed to evaluate their mathematical performance. In this paper, we introduce MathScape, a new benchmark designed to assess the math capabilities of MLLMs using entirely original, leak-free images. Additionally, we propose a novel two-step evaluation method specifically for assessing long answers. MathScape not only challenges existing MLLMs but also aims to inspire the development of more advanced math-focused MLLMs.
Appendix A More Case Study
We present additional case studies. As shown in Figure 10 and Figure 11, the model struggles with complex reasoning tasks. This highlights the need to develop stronger reasoning models.
Figure 10: Illustration of Case Study 2. The red sign marks the incorrect part of the answer.
Figure 11: Illustration of Case Study 3. The red sign marks the incorrect part of the answer.
Appendix B Prompt for Scoring Answers
We summarize the prompt for scoring answers in Figure 12.
Figure 12: Prompt used for scoring answers.
Appendix C Visualization of MathScape
C.1 Visualization of Performance of Different Models.
We include additional math samples in MathScape, as illustrated in Figure 13.
Appendix D Reproducibility Checklist
Question1: Includes a conceptual outline and/or pseudocode description of AI methods introduced?
Answer: Yes.
Question2: Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results?
Answer: Yes.
Question3: Provides well marked pedagogical references for less-familiare readers to gain background necessary to replicate the paper?
Answer: Yes.
Question4: Does this paper make theoretical contributions?
Answer: Yes.
Question5: All assumptions and restrictions are stated clearly and formally?
Answer: Yes.
Question6: All novel claims are stated formally (e.g., in theorem statements)?
Answer: Yes.
Question7: Proofs of all novel claims are included?
Answer: Yes.
Question8: Proof sketches or intuitions are given for complex and/or novel results?
Answer: Yes.
Question9: Appropriate citations to theoretical tools used are given?
Answer: Yes.
Question10: All theoretical claims are demonstrated empirically to hold?
Answer: Yes.
Question11: All experimental code used to eliminate or disprove claims is included?
Answer: Yes.
Question12: Does this paper rely on one or more datasets?
Answer: Yes.
Question13: A motivation is given for why the experiments are conducted on the selected datasets?
Answer: Yes.
Question14: All novel datasets introduced in this paper are included in a data appendix?
Answer: Yes.
Question15: All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes?
Answer: Yes.
Question16: All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations?
Answer: Yes.
Question17: All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available?
Answer: Yes.
Question18: All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing?
Answer: Yes.
Question19: Does this paper include computational experiments?
Answer: Yes.
Question20: Any code required for pre-processing data is included in the appendix?
Answer: Yes.
Question21: All source code required for conducting and analyzing the experiments is included in a code appendix?
Answer: Yes.
Question22: All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes?
Answer: Yes.
Question23: All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from?
Answer: Yes.
Question24: If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results?
Answer: Yes.
Question25: This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks?
Answer: Yes.
Question26: This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics?
Answer: Yes.
Question27: This paper states the number of algorithm runs used to compute each reported result?
Answer: Yes.
Question28: Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information?
Answer: Yes.
Question29: The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank)?
Answer: Yes.
Question30: This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments?
Answer: Yes.
Question31: This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting?
Answer: Yes.
Figure 13: Math problem samples in MathScape.